Let vector B = 5.45 m at 60°. Let vector C have the same magnitude as vector A and a direction angle greater than that of vector A by 25°. Let vector A · vector B = 31.8 m2 and vector B · vector C = 31.5 m2. Find the magnitude and direction of vector A . magnitude m and direction °
Answer #1

Suppose the angle between A and B be X

Angle between A and C be X - 25

A.B = 5.45 p cos(x) = 31.8 ....... (1)

B.C = 5.45 p cos(x - 25) = 31.5 ...... (2)

Dividing 1 and 2

Cos(x - 25)/cos25 = 0.99

(cosx cos25 + sinx sin25)/ cos x = 0.99

Cos25 + sin25 tanx = 0.99

Tan(x) = 0.199

X = 11.28 degree

P = 31.8/5.45 cos(11.28) = 5.95 m

Angle = 60 - 11.28 = 48.72 degree

Comment in case any doubt please rate my answer ...